OpenStudy (anonymous):

A mole of X reacts at a constant pressure of 43.0 atm via the reaction X(g)+4Y(g)→2Z(g), ΔH∘=−75.0 kJ Before the reaction, the volume of the gaseous mixture was 5.00 L. After the reaction, the volume was 2.00 L. Calculate the value of the total energy change, ΔE, in kilojoules.

OpenStudy (aaronq):

Isobaric transition, first law: \(\sf H= \Delta U+w\) for a gas expansion: \(\sf w=P_{ext}*\Delta V\) to convert to joules, you need the gas constants. R = 0.08206 L atm/mol*K, R=8.314 J/mol*K \(\sf w=P_{ext}*\Delta V*\dfrac{8.314 ~J/mol*K}{ 0.08206~L~ atm/mol*K}\) \(\sf \Delta U=\Delta H -[P_{ext}*\Delta V*\dfrac{8.314 ~J/mol*K}{ 0.08206~L~ atm/mol*K}]\) \(\sf \Delta U=−75000~J -[(43.0 atm)*(2-5)L*\dfrac{8.314 ~J}{ 0.08206~L~ atm}]\) Then you need to convert to kJ. by the way U=E, internal energy. More one this: http://en.wikipedia.org/wiki/Isobaric_process